OpenStudy (anonymous):

Is the term (x + 1) a factor of the polynomial shown below? f(x)=x^3-10x^2+27x-12 A. Yes, the remainder is -50. B. No, but the remainder is 0. C. Yes, the remainder is 0. D. No, the remainder is -50.

OpenStudy (anonymous):

when i plugged one into the equation i got 6..

OpenStudy (anonymous):

-50

OpenStudy (anonymous):

D

OpenStudy (anonymous):

oh i see i thought it waqs positive 1 not negative thanks!

OpenStudy (anonymous):

x+1=0 x=-1 f(x)=x^3-10x^2+27x-12 f(x)=(-1)^3-10(-1)^2+27(-1)-12 =-1-10-27-12=-50

OpenStudy (anonymous):

The last one.

OpenStudy (anonymous):

and its not a factor. if you do the polinomyal roots you'll get: (x-4) (x^2-6 x+3) so there's no (x+1) to anyone xD

OpenStudy (anonymous):

meda plz

OpenStudy (anonymous):

it is not a factor and has a remainder -50